OpenStudy (anonymous):

Ok for all you smart people out there.. See if you can figure this out! if a object with a mass of 4kg moving at 7.5m/s in the same direction as a 5kg object moving at 2m/s collide with another, and continue to move in the same direction. what is the speed of the 5kg object if the 4kg object continues to move with a speed of 3m/s?

OpenStudy (roadjester):

This would be an inelastic collision but not perfectly inelastic.

OpenStudy (roadjester):

\( \huge m_1{v_1}_i +m_2{v_2}_I=m_1{v_1}_f +m_2{v_2}_f \)

OpenStudy (anonymous):

momentum in = momentum out each ball has it's own momentum which is the "in" momentum. The "out" momentum is what happens after the collision (mass1 * velocity1) + (mass2 * velocity2) = (mass1 * velocity1)final + (mass2 * velocity2)final plug in the numbers: (4kg * 7.5 m/s) + (5kg * 2 m/s) = (4kg * 3 m/s) + ( 5kg * X m/s) solve the equation for X: 40 kg m/s = 12 kg m/s + 5X kg m/s 28 kg m/s = 5X kg m/s X = 5.6 m/s velocity............ for the 5kg object after impact

OpenStudy (anonymous):

you are right @biire

OpenStudy (anonymous):

Thanks you guys :) I don't understand half the stuff we do in ICP haha